OpenStudy (anonymous):

-3(4x+3) +4(6x+1) = 43 -12x-9 + 24x+4 =43 +9 +9 -12x +24 +13=43 i still dont know where im going wrong

HanAkoSolo (jamierox4ev3r):

-3(4x+3)+4(6x+1)=43 This is the equation, right? @datgrlyouenvy ?

OpenStudy (anonymous):

correct. i need to know where i went wrong and help finishing

OpenStudy (anonymous):

OK, why did you add 9 twice to that one side?

OpenStudy (goldphenoix):

-12x-9 + 24x+4 =43 This is an equation, you must simplify the left side.

OpenStudy (anonymous):

I suggest you start from: -12x-9+24x+4=43 and combine like terms.

OpenStudy (goldphenoix):

You cannot add 9 two times on one side of th equation.

OpenStudy (goldphenoix):

On the left side, which terms can we combine?

OpenStudy (anonymous):

oh crap i just noticed that. i meant to do it under 43 my dislexia toys with me a lot when im doing math

OpenStudy (anonymous):

You may find it helpful to use graph paper to help keep your numbers in line.

OpenStudy (anonymous):

we can add the -9 and 4 together. to make -5 correct?

OpenStudy (goldphenoix):

Yep! What other terms can we combine?

OpenStudy (anonymous):

-12x and 24x to make 36

OpenStudy (goldphenoix):

Hmm. Not quite. The value should be less than 24, since we have -12x. :)

OpenStudy (anonymous):

so you'd subtract 12 from 24 to make 12

OpenStudy (goldphenoix):

Correct! Make sure to put the variable, x. So now our equation reads: \[\large \large 12x - 5 =43\] What do we do next?

OpenStudy (anonymous):

subtract 5 from the -5 and 43 to make 12x=38

OpenStudy (goldphenoix):

Great. What is the final step?

OpenStudy (anonymous):

dividing them both by 12

OpenStudy (goldphenoix):

Hmm. wait.

OpenStudy (goldphenoix):

Yeah, I know. >_<

OpenStudy (anonymous):

what was missed

OpenStudy (goldphenoix):

Oooh. You don't subtract 43 by 5.

OpenStudy (goldphenoix):

How can we get rid of -5?

OpenStudy (anonymous):

i said that above

OpenStudy (goldphenoix):

Well. -5 - 5 = -10. Tha does not cancel out. So what is another possibility?

OpenStudy (anonymous):

thats how i got the 38

OpenStudy (anonymous):

subtracting 43 from its self and subtracting 43 from -5

OpenStudy (goldphenoix):

Hmm. Wait, tell me this. How can -5 turn into 0?

OpenStudy (goldphenoix):

What do you have to do for -5 to = 0 ? If that made sense.

OpenStudy (anonymous):

adding

OpenStudy (goldphenoix):

Add -5 by what number?

OpenStudy (anonymous):

5

OpenStudy (goldphenoix):

Great! So, -5 + 5 = 0 You must do it to the other side. So, 43 + 5 = ?

OpenStudy (anonymous):

48

OpenStudy (goldphenoix):

Great. Now like you said before, divide 12. 48 / 12 = ?

OpenStudy (anonymous):

4

OpenStudy (goldphenoix):

Great! Therefore, x = 4. I hope this help! :)

OpenStudy (anonymous):

a lot thanks. sometimes im really good at equations and others my dislexia gets the better

OpenStudy (goldphenoix):

Not bad. Even though you have dislexia, you are smarter than some people I know who doesn't have it.

OpenStudy (anonymous):

lmao im 29 and just learning. started at fractions about 2 months ago so in actuality its taken many many years to get here and 2 great/patient teachers who know how to do problems multiple ways and not just one